Find all real solutions of the given equation. x+x^1/2-20=0
Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

x=16
check
16+4-20=0
ok
(sqrt(x)-4) (sqrt(x)+5) = 0

Answer by Alan3354(69443)

x+x^1/2-20=0
--------------
The easiest way to do this one is to treat it as a quadratic in x^1/2.
Use xh for x^1/2
(xh)^2 + (xh) - 20 = 0
(xh + 5)*(xh - 4) = 0
xh = -5 --> x = 25
xh = 4 --> x = 16
------------
Much simpler and less error chance than expanding it, etc.